OpenStudy (anonymous):

Is anybody willing to help me with my homework and getting a better understanding of Linear Functions? I'm struggling with math and my teacher isn't much help. I would greatly appreciate it.

OpenStudy (aravindg):

I can try.

OpenStudy (anonymous):

What do you need for help?

OpenStudy (anonymous):

Okay thank you both! And just like everything, I don't understand anything dealing with this subject.

OpenStudy (aravindg):

lol thats a lot. Can you specify some part or a specific question. Because it takes an entire lecture to explain it fully.

OpenStudy (aravindg):

If you need to learn about straight lines, you can check my tutorial: http://openstudy.com/updates/50f8cc3fe4b027eb5d9a556a

OpenStudy (anonymous):

I know lol Uhm function notation and graphs. Heres an example. You are in charge of finding a place to have your little sister’s birthday party. You call around, but many places are already booked. You find one place, Jump and Fun, which looks promising. Here's what you learn when you call Jump and Fun: There is a flat fee of $120 to rent a room. There is an additional charge of $5 per child for access to the bounce houses and miniature golf course. The more friends your sister invites, the more the party will cost. So, how many friends can she invite? You can use an algebraic expression to find out.

OpenStudy (anonymous):

Straight lines? I will take a look at it right now.

OpenStudy (aravindg):

Is the amount you have with you mentioned?

OpenStudy (anonymous):

I have no clue what amount your talking about. Thats just what the lesson was telling me trying to give an example.

OpenStudy (anonymous):

c increases by $25. This means for every 5 kids that attend the party, the cost increases by $25. the equation would be c = 120 + 5x its a function because there are no repeating x values (input values). In other words, if 5 kids attend the party the cost is $145. There is not another cost that is associated with 5 kids attending the party.

OpenStudy (anonymous):

Do you know about the domain and range of this function to? @bullkidd

OpenStudy (aravindg):

In fact the algebraic expression is in fact an inequality \[c \ge 120+5x\] The maximum amount you can pay is the money you have. So c=120+5x Once you know C you can find x where x is max number of kids that can be invited.

OpenStudy (anonymous):

so with that equation, you would have to determine the cost of the party if 10,15,20,25 and 30 kids attended! you would have to create a table of values!! so in the "x" side you will write down those numbers that i just mentioned and the other side you have "c" wich is the total cost of the party.

OpenStudy (anonymous):

Yeah thats right @AravindG ! And thats also right @bullkidd ! You guys are really good at this. This is not my strongest subject at all.

OpenStudy (anonymous):

so the table should look something like this,...(dont forget the equation is c=120+5(x) and all you have to do is add the numbers of the kids that might attend to 'x' which for the first one would be c=120+5(10) -> c=120+50 -> c=170$ so you would have to do that for each and i wrote the numbers earlier which would be 10,15,20,25, and 30). x | c 10 | 170 15 | 195 20 | 220 25 | 245 30 | 270 so because your budget is 250$ you would decide to let 25 kids in, because that is less than your budget and its affordable =\ i hope that helped!!

OpenStudy (anonymous):

It did! But do you know how to get the domain and range of this function? That was also something apart of it @bullkidd

OpenStudy (aravindg):

To explain the concept of domain and range we can take the same example. Assume the max number of kids that can be invited be y. Then domain of x is set of all possible values x can take. So here it becomes 0,1,2,3...y 0 means no friends invited,1 mean 1 friend invited y means max number of friends according to budget invited. We can write this in notation [0,y] where closed bracket means it includes all values of x from 0 to y including 0 and y. Now once u understand domain its simple to understand range. Range is the set of values the output can take while domain was set of all input values ie no of kids. Here output is the total cost that you have to pat. Assume the money you have is K. Then c<=K. So range is [120.,K]. Why start from 120? Because that is the c value when 0 input given. When max input given ie x=y we get c=K. So range can be written as [120,K].

OpenStudy (anonymous):

Domain would be all the 'x' values of the problem which is 10,15,20,25 and 30. And the Range would be all the 'y'(in this case 'C') values of the problem which would be 170, 195, 220, 245 and 270!!

OpenStudy (anonymous):

Wow thank you both! You guys are awesome(: If I have anymore trouble (which I will lol sorry) would you guys mind helping me out another time? @AravindG and @bullkidd

OpenStudy (aravindg):

Indeed! Good luck on your studies! :)

OpenStudy (anonymous):

Definitely ^_^ and good luck with the teachers :S lol =)